1. Marigold Corp. estimates its annual warranty expense as 3% of annual net sales. The following data relate to the calendar year 2021:

Net sales $1506000
Warranty liability account
    Balance, Dec. 31, 2020 $13000 debit before adjustment
    Balance, Dec. 31, 2021 32180 credit after adjustment


Which one of the following entries was made to record the 2021 estimated warranty expense?

a

Warranty Expense 32180
    Warranty Liability 32180

b

Warranty Expense 37635
Retained Earnings (prior-period adjustment) 7545
    Warranty Liability 45180

c

Warranty Expense 45180
    Warranty Liability 45180

d

Warranty Expense 45180
    Retained Earnings (prior-period adjustment) 7545
    Warranty Liability 37635

2. Crane Company has 60 employees who work 8-hour days and are paid hourly. On January 1, 2020, the company began a program of granting its employees 10 days of paid vacation each year. Vacation days earned in 2020 may first be taken on January 1, 2021. Information relative to these employees is as follows:

Year Hourly
Wages
Vacation Days Earned
by Each Employee
Vacation Days Used
by Each Employee
2020 $26.50 10 0
2021 28.50 10 8
2022 31.50 10 10


Crane has chosen to accrue the liability for compensated absences at the current rates of pay in effect when the compensated time is earned.

What is the amount of the accrued liability for compensated absences that should be reported at December 31, 2022?

a

$203760.

b

$405360.

c

$151200.

d

$178560.

3. Vaughn Co. is a retail store operating in a state with a 7% retail sales tax. The retailer may keep 2% of the sales tax collected. Vaughn Co. records the sales tax in the Sales Revenue account. The amount recorded in the Sales Revenue account during May was $762910.

The amount of sales taxes (to the nearest dollar) for May is

a

$62846.

b

$49910.

c

$53405.

d

$68453
Answer #1

Bc E A 1 2 Warranty Expense 3 Warranty Liability 45180 45180 2 2021 27360 2022 151200 9 Total 178560 10 11 = Hourly x 8 hours11 2 Warranty Expense 3 Warranty Liability =1506000*3% =+B2 52 7 2021 =28.5*8*2*60 8 2022 =31.5*8*10*60 9 Total =SUM(B7:38) 1